OpenStudy (anonymous):

Solve: 2x^3 + x^2 + 5x - 3 = 0 Step by step anyone? Much appreciated! :)

OpenStudy (anonymous):

\[2^{3} + x^{2} +5x - 3 = 0\] \[8 + x^{2} +5x - 3 = 0\] \[5 + x^{2} +5x = 0\] can you get X alone?

OpenStudy (anonymous):

i think, there are no solution there

OpenStudy (anonymous):

x = -1 or -5...? >.>

OpenStudy (anonymous):

it's a trinomial so it comes to be x(x + 5) = -5

OpenStudy (anonymous):

if x=-1, ---> 2(-1)^3+(-1)^2+5(-1)-3=-2+1-5-3=-9 not equal zero if x=-5, ---> 2(-5)^3+(-5)^2+5(-5)-3=-250+25-25-3=-253 not equal zero, so x=-1 or x=-5 is not solution

OpenStudy (anonymous):

@.@... I'm so confused now....

OpenStudy (anonymous):

it's a no solution